# Sabiá-3 Technical Report

### v1

@@ Abstract
This report presents Sabiá-3, our new flagship language model trained on a large brazilian-centric corpus. Evaluations across diverse professional and academic benchmarks show a strong performance on Portuguese and Brazil-related tasks. Sabiá-3 shows large improvements in comparison to our previous best of model, Sabia-2 Medium, especially in reasoning-intensive tasks. Notably, Sabiá-3’s average performance matches frontier LLMs, while it is offered at a three to four times lower cost per token, reinforcing the benefits of domain specialization.
@@ 1 Introduction
This technical report presents the details of the development and evaluation of the Sabiá-3 model. We trained Sabiá-3 on a large corpus of documents written in Portuguese, with a special focus on Brazil-related resources. Through its training, the model was exposed to information relevant to the Brazilian culture, history, and context. The main objective was to have a specialized model that is aware of the linguistic nuances, societal norms, and regional variations unique to the country. Throughout this report, we show that this specialization allows the model to perform better in knowledge-intensive tasks.
We applied an approach of continual learning by leveraging a “generalist” model that already acquired some level of language understanding and reasoning abilities, and then further train it on our corpus of high-quality data relevant to the Brazilian context. The development consisted of two main phases: (1) the pre-training phase, in which we further train a pre-trained model on specialized data following a self-supervised learning strategy optimizing for the next token prediction objective, and (2) the post-training phase where the model is tuned to follow instructions and align to human preferences.
@@
Figure 1: Price in USD per million tokens, considering an equal proportion of input and output tokens, versus performance on 52 Brazilian exams (in Portuguese). The dashed curve represents a Pareto frontier for general-purpose LLMs such as Llama 3.1 and GPT-4o, which Sabiá-3 surpasses due to its domain specialization.
@@ 2 Evaluation
In this section, we compare Sabiá-3 with proprietary and open-source LLMs, listed in Table 1, on various benchmarks.
@@ 2.1 Multiple-choice Exams
We have developed a benchmark suite to evaluate the Sabiá-3 performance on 78 academic exams, including entry-level, undergraduate, and professional certification exams. Our focus is on multiple-choice questions from various Brazilian educational assessments, particularly those administered after the middle of 2023 to mitigate the risk of data contamination. However, ENADE is an exception, including spans of 2022 and 2023 to cover a wide range of disciplines.
@@
Table 2 shows the scores of the models in ten benchmarks derived from academic and professional exams applied after mid-2023. The results illustrate that Sabiá-3 has a comparable level of accuracy to GPT-4o across all 52 exams assessed. GPT-4o exhibits an advantage particularly in the Medicine benchmarks, while Sabiá-3 demonstrates a notable performance in the CPNU exams. Furthermore, Sabiá-3 surpasses Llama-3.1 405B by 3 percentage points in accuracy, performing equally or better in 63% of the exams. Claude 3.5 Sonnet achieves the highest average accuracy, although it presents a higher risk of data contamination due to its training data cutoff date.
Llama-3.1 Llama-3.1 GPT-4o GPT-4o Claude 3.5 Sabiá-2 Sabiá-3
@@
ENEM Multiple 1 84.4% 85.5% 84.9% 84.9% 92.7% 77.1% 86.6%
BLUEX Multiple 2 77.3% 83.9% 85.1% 88.7% 93.0% 75.9% 86.2%
ENADE Multiple 28 66.2% 73.0% 63.2% 75.2% 78.0% 64.6% 77.2%
CPNU Multiple 8 82.2% 88.9% 83.4% 88.5% 83.8% 79.2% 90.6%
POSCOMP Comp. Sci 1 61.5% 78.5% 63.1% 73.8% 76.9% 50.8% 76.9%
OAB Law 4 69.3% 77.4% 63.3% 83.1% 80.6% 72.1% 75.9%
ENAM Law 2 50.6% 56.2% 50.6% 66.2% 75.0% 53.1% 69.4%
CFCES Account. 1 50.0% 60.0% 60.0% 74.0% 76.0% 64.0% 76.0%
Revalida Medicine 3 81.3% 84.2% 74.8% 87.1% 85.7% 73.0% 83.5%
MREX Medicine 2 61.0% 61.5% 56.6% 68.7% 73.4% 49.9% 57.2%
Average - 52 69.4% 75.9% 67.4% 78.6% 80.0% 67.3% 78.9%
Table 2: Accuracy on academic benchmarks. The average results (last column) cover all the 52 exams applied after the middle of 2023, which is the knowledge cutoff date of Sabiá-3. Claude 3.5’s cutoff date is April of 2024, so exams released before that date might be present in its training data.
In Figure 1, we show the cost per token as of October 2024 in comparison to the average accuracy on the aforementioned exams. We have plotted a Pareto curve based on generalist models, i.e., the ones not specialized in the Brazilian context. Sabiá-3 stays above this frontier, underscoring the cost-quality benefits that result from domain specialization.
@@ 2.3 Instruction-Following Capabilities
The results presented in Table 3 show that Sabiá-3 outperforms Sabiá-2 Medium in terms of instruction-following capabilities. However, it lags behind other models evaluated in this study, including cheaper options such as GPT-4o Mini and Llama-3.1 8B.
@@ 2.4 Long Context Capabilities
Sabiá-3 can process up to 32,000 tokens in a sequence. Here we evaluate how well the model uses this capacity through the Needle-in-the-Haystack (NIAH) benchmark [9], which measures the model’s ability to find answers within a long, unrelated context. The test involves inserting a small piece of text within a larger context ranging from 1,000 to 32,000 tokens, at various depths from 0% to 100%. Since we are focusing on evaluating the performance of an LLM on Portuguese tasks, we adapt this benchmark using the book Dom Casmurro [^1] https://machadodeassis.net/texto/dom-casmurro/11503 as the context. In a random spot within the text, we insert as a needle a sentence that says “O número mágico de Campinas é {random_number}.” (“The magic number of Campinas is {random_number}.”), replacing {random_number} with a 6-digit number. Then, we provide this modified context along with the question “Qual o número mágico de Campinas?” (“What is the magic number of Campinas?”) to the model. Using regular expression, we check whether the model correctly outputs the random number as the answer.
As depicted in Figure 5, Sabiá-3 exhibited a perfect recall in our Portuguese version of the NIAH benchmark, consistently locating relevant information within long contexts. We acknowledge, however, that the task, due to its simplicity, may not fully capture the long-context capabilities of language models. Therefore, for a better measure of long-range comprehension skills, we need Portuguese versions of more challenging benchmarks, such as the QuALITY benchmark [14], which tests a model’s ability to answer questions about books.
@@ 2.5 Function Calling Capabilities
To evaluate the models we use the original examples in English. Table 4 shows the results of Sabiá-3 and other proprietary LLMs. The results indicate that Sabiá-3 ranks competitively alongside top-performing models. It outperforms in the Non-live categories but requires enhancements in multi-turn and multi-step tasks. For instance, increasing the support for more than 32,000 tokens could improve in one of the multi-turn categories. It also needs to improve in the irrelevance category.
@@ 3 Conclusion
We introduced Sabiá-3, a language model focused on Brazil-related content. It improves upon our earlier Sabiá-2 models, especially in handling long texts, reasoning, and coding. Despite rivaling state-of-the-art proprietary and open-source models on knowledge-intensive tasks while having a lower per token cost, Sabiá-3 still has room for improvement in multi-step tasks and following instructions accurately, both of which we plan to address in future work.

### v2

@@ Abstract
This report presents Sabiá-3, our new flagship language model, and Sabiazinho-3, a more cost-effective sibling. The models were trained on a large brazilian-centric corpus. Evaluations across diverse professional and academic benchmarks show a strong performance on Portuguese and Brazil-related tasks. Sabiá-3 shows large improvements in comparison to our previous best of model, Sabia-2 Medium, especially in reasoning-intensive tasks. Notably, Sabiá-3’s average performance matches frontier LLMs, while it is offered at a three to four times lower cost per token, reinforcing the benefits of domain specialization.
@@ 1 Introduction
This technical report presents the details of the development and evaluation of the Sabiá-3 and Sabiazinho-3 models. We trained them on a large corpus of documents written in Portuguese, with a special focus on Brazil-related resources. Through training, models were exposed to information relevant to Brazilian culture, history, and context. The main objective was to have a specialized model that is aware of the linguistic nuances, societal norms, and regional variations unique to the country. Throughout this report, we show that this specialization allows the models to perform better in knowledge-intensive tasks.
We applied an approach of continual learning by leveraging a “generalist” model that already acquired some level of language understanding and reasoning abilities, and then further trained it on our corpus of high-quality data relevant to the Brazilian context. The development consisted of two main phases: (1) the pre-training phase, in which we further train a pre-trained model on specialized data following a self-supervised learning strategy optimizing for the next token prediction objective, and (2) the post-training phase where the model is tuned to follow instructions and align to human preferences.
@@
Figure 1: Price in USD per million tokens, considering an equal proportion of input and output tokens, versus performance on 70 Brazilian exams (in Portuguese). The dashed curve represents a Pareto frontier for general-purpose LLMs such as Llama 3.1 and GPT-4o, which Sabiá-3 and Sabiazinho-3 surpass due to their domain specialization.
@@ 2 Evaluation
In this section, we compare Sabiá-3 and Sabiazinho-3 with proprietary and open-source LLMs, listed in Table 1, on various benchmarks.
@@ 2.1 Multiple-choice Exams
We have developed a benchmark suite to evaluate the Sabiá-3 performance on 96 academic exams, including entry-level, undergraduate, and professional certification exams. Our focus is on multiple-choice questions from various Brazilian educational assessments, particularly those administered after the middle of 2023 to mitigate the risk of data contamination. However, ENADE is an exception, including spans of 2022 and 2023 to cover a wide range of disciplines.
@@
Table 2 shows the scores of the models in ten benchmarks derived from academic and professional exams applied after mid-2023. The results illustrate that Sabiá-3 has a comparable level of accuracy to GPT-4o across all 70 exams assessed. GPT-4o exhibits an advantage particularly in the Medicine benchmarks, while Sabiá-3 demonstrates a notable performance in the CPNU exams. Furthermore, Sabiá-3 surpasses Llama-3.1 405B by 3 percentage points in accuracy, performing equally or better in 63% of the exams. Claude 3.5 Sonnet achieves the highest average accuracy, although it presents a higher risk of data contamination due to its training data cutoff date.
Llama-3.1 Llama-3.1 GPT-4o GPT-4o Claude 3.5 Sabiá-2 Sabiazinho-3 Sabiá-3
@@
ENEM Multiple 2 83.8% 84.9% 83.8% 87.2% 89.9% 71.8% 81.0% 87.7%
BLUEX Multiple 2 77.3% 83.9% 85.1% 88.7% 93.0% 75.9% 79.5% 86.2%
ENADE Multiple 28 66.2% 73.0% 63.2% 75.2% 78.0% 64.6% 69.2% 77.2%
CPNU Multiple 8 82.2% 88.9% 83.4% 88.5% 83.8% 79.2% 84.4% 90.6%
BNDES Multiple 14 76.5% 80.0% 76.5% 86.7% 81.2% 75.5% 77.8% 84.3%
BCB Multiple 3 75.8% 79.4% 76.3% 84.1% 81.8% 65.0% 73.9% 80.1%
POSCOMP Comp. Sci 1 61.5% 78.5% 63.1% 73.8% 76.9% 50.8% 56.9% 76.9%
OAB Law 4 69.3% 77.4% 63.3% 83.1% 80.6% 72.1% 69.9% 75.9%
ENAM Law 2 50.6% 56.2% 50.6% 66.2% 75.0% 53.1% 53.1% 69.4%
CFCES Account. 1 50.0% 60.0% 60.0% 74.0% 76.0% 64.0% 54.0% 76.0%
Revalida Medicine 3 81.3% 84.2% 74.8% 87.1% 85.7% 73.0% 70.9% 83.5%
MREX Medicine 2 61.0% 61.5% 56.6% 68.7% 73.4% 49.9% 47.9% 57.2%
Average - 70 71.3% 77.0% 69.9% 80.6% 80.5% 68.9% 72.1% 80.2%
Table 2: Accuracy on academic benchmarks. The average results (last row) cover all the 70 exams applied after the middle of 2023, which is the knowledge cutoff date of Sabiá-3. Claude 3.5’s cutoff date is April of 2024, so exams released before that date might be present in its training data.
In Figure 1, we show the cost per token as of October 2024 in comparison to the average accuracy on the aforementioned exams. We have plotted a Pareto curve based on generalist models, i.e., the ones not specialized in the Brazilian context. Sabiá-3 and Sabiazinho-3 stay above this frontier, underscoring the cost-quality benefits that result from domain specialization.
@@ 2.3 Instruction-Following Capabilities
The results presented in Table 3 show that Sabiá-3 and Sabiazinho-3 outperform Sabiá-2 Medium in terms of instruction-following capabilities. However, it lags behind other models evaluated in this study, including cheaper options such as GPT-4o Mini and Llama-3.1 8B.
@@ 2.4 Long Context Capabilities
Sabiá-3 and Sabiazinho-3 can process up to 32,000 tokens in a sequence. Here we evaluate how well the model uses this capacity through the Needle-in-the-Haystack (NIAH) benchmark [9], which measures the model’s ability to find answers within a long, unrelated context. The test involves inserting a small piece of text within a larger context ranging from 1,000 to 32,000 tokens, at various depths from 0% to 100%. Since we are focusing on evaluating the performance of an LLM on Portuguese tasks, we adapt this benchmark using the book Dom Casmurro [^1] https://machadodeassis.net/texto/dom-casmurro/11503 as the context. In a random spot within the text, we insert as a needle a sentence that says “O número mágico de Campinas é {random_number}.” (“The magic number of Campinas is {random_number}.”), replacing {random_number} with a 6-digit number. Then, we provide this modified context along with the question “Qual o número mágico de Campinas?” (“What is the magic number of Campinas?”) to the model. Using regular expression, we check whether the model correctly outputs the random number as the answer.
As depicted in Figures 5 and 6, Sabiá-3 exhibited a perfect recall in our Portuguese version of the NIAH benchmark and Sabiazinho-3 reached a recall of 99.48%, consistently locating relevant information within long contexts. We acknowledge, however, that the task, due to its simplicity, may not fully capture the long-context capabilities of language models. Therefore, for a better measure of long-range comprehension skills, we need Portuguese versions of more challenging benchmarks, such as the QuALITY benchmark [14], which tests a model’s ability to answer questions about books.
@@
Figure 6: Performance of the Sabiazinho-3 model in the Portuguese-adapted Needle-in-the-Haystack (NIAH) benchmark.
@@ 2.5 Function Calling Capabilities
To evaluate the models we use the original examples in English. Table 4 shows the results of Sabiá-3, Sabiazinho-3, and other proprietary LLMs. The results indicate that Sabiá-3 ranks competitively alongside top-performing models. It outperforms in the Non-live categories but requires enhancements in multi-turn and multi-step tasks. For instance, increasing the support for more than 32,000 tokens could improve in one of the multi-turn categories. It also needs to improve in the irrelevance category.
@@ 3 Conclusion
We introduced Sabiá-3 and Sabiazinho-3, a family of language models focused on Brazil-related content. They improve upon our earlier Sabiá-2 models, especially in handling long texts, reasoning, and coding. Despite rivaling state-of-the-art proprietary and open-source models on knowledge-intensive tasks while having a lower per token cost, Sabiá-3 and Sabiazinho-3 still have room for improvement in multi-step tasks and following instructions accurately, both of which we plan to address in future work.